In order to conduct a hypothesis test for the population proportion, you sample 450 observations that result in 189 successes. (You may find it useful to reference the appropriate table: z table or t table)

H0: p ≥ 0.45; HA: p < 0.45.

a-1. Calculate the value of the test statistic. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

TEST STATISTIC =

a-2. Find the p-value.

  • 0.025   p-value < 0.05
  • 0.05   p-value < 0.10
  • p-value   0.10
  • p-value < 0.01

  • 0.01   p-value < 0.025

a-3. At the 0.10 significance level, What is the conclusion?

  • Do not reject H0 since the p-value is smaller than significance level.

  • Do not reject H0 since the p-value is greater than significance level.

  • Reject H0 since the p-value is smaller than significance level.

  • Reject H0 since the p-value is greater than significance level.

a-4. Interpret the results at αα = 0.10

  • We cannot conclude that the population mean is less than 0.45.

  • We conclude that the population mean is less than 0.45.

  • We cannot conclude that the population proportion is less than 0.45.

  • We conclude that the population proportion is less than 0.45.

H0: p = 0.45; HA: p ≠ 0.45.

b-1. Calculate the value of the test statistic. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

TEST STATISTIC =

b-2. Find the p-value.

  • 0.01   p-value < 0.025
  • 0.025   p-value < 0.05
  • 0.05   p-value < 0.10
  • p-value   0.10
  • p-value < 0.01

b-3. At the 0.10 significance level, What is the conclusion?

  • Reject H0 since the p-value is greater than significance level.

  • Reject H0 since the p-value is smaller than significance level.

  • Do not reject H0 since the p-value is greater than significance level.

  • Do not reject H0 since the p-value is smaller than significance level.



b-4. Interpret the results at αα = 0.10.

  • We conclude that the population mean differs from 0.45.

  • We cannot conclude that the population mean differs from 0.45.

  • We conclude the population proportion differs from 0.45.

  • We cannot conclude that the population proportion differs from 0.45.
Answer #1

Part a1)

P = X / n = 189/450 = 0.42

Test Statistic :-
Z = ( P - P0) / √(P0 * q0 / n)
Z = ( 0.42 - 0.45 ) / √(( 0.45 * 0.55) /450)
Z = -1.28

Part a2)

P value = P ( Z < -1.28 ) = 0.1004
Looking for the value Z = -1.28 in standard normal table to find the P value

p-value    0.10

Part a3)

Reject null hypothesis if P value < α = 0.1
Since P value = 0.1004 > 0.1, hence we fail to reject the null hypothesis
Conclusion :- We Fail to Reject H0

Do not reject H0 since the p-value is greater than significance level.

Part a4)

We cannot conclude that the population proportion is less than 0.45.

Part b1)

P = X / n = 189/450 = 0.42

Test Statistic :-
Z = ( P - P0) / √(P0 * q0 / n)
Z = ( 0.42 - 0.45 ) / √(( 0.45 * 0.55) /450)
Z = -1.28

Part b2)

P value = 2 * P ( Z < -1.28 ) = 0.2008
Looking for the value Z = -1.28 in standard normal table to find the P value

p-value    0.10

Part b3)

Reject null hypothesis if P value < α = 0.1
Since P value = 0.2008 > 0.1, hence we fail to reject the null hypothesis
Conclusion :- We Fail to Reject H0

Do not reject H0 since the p-value is greater than significance level.

Part b4)

We cannot conclude that the population proportion differs from 0.45.